OpenStudy (anonymous):

The current I in an electrical conductor varies inversely as the resistance R of the conductor. The current is 2 amperes when the resistance is 990 ohms. What is the current when the resistance is 570 ohms?

OpenStudy (anonymous):

divide 990 by two.

OpenStudy (amistre64):

no replies yet eh.... its like throwing those rings on the bottle necks i tell ya....

OpenStudy (anonymous):

take that value and subtract it from 570. find percentage etcetcetc

OpenStudy (anonymous):

lol :P

OpenStudy (anonymous):

im helping someone else out amistre take over :P

OpenStudy (amistre64):

2 = k/990 k = 1980 y = 1980/570 y = 3.47...

OpenStudy (anonymous):

......

OpenStudy (amistre64):

there should be jeopardy music playing in the backgrond

OpenStudy (anonymous):

it worked! ....again.!

OpenStudy (amistre64):

*shock*..... i got lucky lol

OpenStudy (anonymous):

lol, now she is really ticked! she is gonna post one that is bothering her.....